# A Novel $J_{M}$-Based Method for Interpolation in Metric Spaces

### v1

@@ Abstract
We present a new interpolation method for metric spaces, termed the $J_{M}^{\prime}$-interpolated metric space. Building on previous work where two interpolation methods were developed in analogy with Peetre’s classical $K$ and $J$ methods, this approach introduces a distinct construction based on approximating sequences in the intersection space $X_{0}\cap X_{1}$ equipped with the metric $J_{M}(1)$. The method defines a relative completion within the union space $X_{0}\cup X_{1}$, yielding a well-defined interpolated metric space with controllable relationships between the $J_{M}$ and $K_{M}$ metrics. While we do not compare this method to the previously developed approaches nor explore its correspondence with classical interpolation functors in normed spaces, the construction enriches the set of tools available for metric-space interpolation theory. Furthermore, as expected, these methods still preserve the Lipschitz property of an operator between two interpolated spaces. Potential applications include general Nonlinear Analysis and Wasserstein-type spaces where the sequence-based nature of the method may provide practical advantages.
@@ Introduction
In a previous work published in 2024 [see [13]], we introduced two methods of interpolation for general metric spaces, inspired by the classical $K$ and $J$ methods in the normed setting of Peetre [see [10] and [3]]. These constructions allowed the extension of classical interpolation ideas to the broader context of metric spaces, defining analogues of the $K$- and $J$-interpolation spaces and establishing a structured approach to study interpolation without relying on linearity or normed assumptions. This type of interpolation acts on metric spaces and preserves the Lipschitz continuity of operators in the interpolated spaces, in a manner analogous to interpolation functors in classical Banach space theory. The methods provided a framework for constructing interpolated metric spaces that capture intermediate structures between two given metric spaces, opening avenues for applications in areas where metric spaces arise naturally, such as nonlinear analysis, optimal transport, and metric geometry.
In this article, we develop a new interpolation method, the $J_{M}^{\prime}$-interpolated metric space, which complements the previously introduced methods by offering a fundamentally different construction. Unlike the prior methods, which relied on direct adaptations of the $K$ and $J$ functionals, this new method is based on approximating sequences within the intersection space $X_{0}\cap X_{1}$, endowed with the metric $J_{M}(1)$, and its relative completion inside the union space $X_{0}\cup X_{1}$. This sequence-based approach introduces a unique perspective, which could prove advantageous in situations where the behavior of sequences in the intersection space is critical or where the structure of the union space provides natural constraints.
The primary motivation behind proposing a new interpolation method is that expanding the toolbox of available interpolation techniques enriches the study of metric-space interpolation, allowing researchers to choose the most suitable method depending on the specific problem or application. This methodological diversity is particularly relevant for applications to Nonlinear Analysis and spaces with nontrivial geometries, such as Wasserstein-type spaces [see [1] and [15] and [16]], where the sequence-oriented nature of this approach might yield novel insights or simplifications.
It is important to note that the goal of this article is the construction of the new interpolation method itself. We do not investigate its relationship with the previously developed methods, nor do we attempt to determine whether it generates interpolation functors in normed spaces that are equivalent to classical constructions by Peetre or Lions. Our focus is on establishing the theoretical foundation of this new approach and analyzing its intrinsic properties, such as its metric structure, continuity of inclusions, and basic interpolation inequalities. As a final remark, we suggest that future research could explore potential applications of this method, particularly in Wasserstein-type spaces, where the specific sequence-based construction might reveal previously inaccessible properties of interpolated structures.
@@ 1.1 Relative Completion of Metric Spaces
We define the relative completion of $(A,d_{A})$ in $(B,d_{B})$, denoted by $A_{B}$, by
$A_{B}:=\{x\in B;\,\,\textbf{$\exists$ Cauchy sequence $(x_{n})$ in $(A,d_{A})$ and $d_{B}(x_{n},x)\to 0$}\}.$
@@ 1.2 Pairs of Compatible Metric Spaces
Let $(X_{0},X_{1})_{X}$ and $t>0$. For $(x,y)\in X_{0}^{2}\cup X_{1}^{2}$, define the function $h_{t}$ which will be given by
@@ 1.3 The $K_{M}$ Functional and its properties
$K_{M}(a;x,y)\leq\max\biggr\{1,\frac{a}{b}\biggr\}\,K_{M}(b;x,y)$ (3)
@@ 3 The $J_{M}^{\prime}$-Interpolated Metric Space
Now, using Lemma 1.5, with $a=1$ and $b=2^{j}$ in each term $K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y})$, we have
@@
Definition 3.4 (The $J_{M}^{\prime}$-interpolated metric space).
The $J_{M}^{\prime}$-interpolated metric space is denoted by $(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$ and is the relative completion of the metric space
@@
Remark 3.2.
An important fact is that the interpolated spaces above were defined as the completion of the intersection $X_{0}\cap X_{1}$ with a specific metric in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. This ensures that the elements of the interpolated spaces are contained in $X_{0}\cup X_{1}$. However, note that such metrics on the intersection also satisfy the conditions for completion in $(X,d_{X})$ which can be substantially “larger” than $X_{0}\cup X_{1}$, providing many more new elements to the completions. That is why the name ”relative” was chosen for this type of completion. We do not study here the advantages of completing the spaces beyond $X_{0}\cup X_{1}$.
4 Properties of the $J_{M}^{\prime}$-Interpolated Metric Space
@@
The inclusions $X_{0}\cap X_{1}\subseteq(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}\subseteq X_{0}\cup X_{1}$ are valid and furthermore
@@
    $K_{M}(1;x,y)\leq M_{\theta,q}\cdot\Delta_{\theta,q}(x,y),\quad x,y\in(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$.
@@
Lemma 4.2 (Reindexing and interpolation factor).
@@
Then, $T_{\theta,q}:=T|_{(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}}:(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}\to(\overrightarrow{Y}_{\theta,q}^{J_{M}})^{\prime}$ is Lipschitzian with Lipschitz constant not exceeding $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}>0$.
@@
where in the last inequality we used the previous Lemma 4.2. From this inequality, we have that
@@
$\Delta_{\theta,q;Y}(T(x),T(y))\leq 2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$ (19)
@@
$\Delta_{\theta,q;Y}(T(x),T(y))\leq\max\{\omega_{0},\omega_{1}\}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}.$ (20)
@@
Remark 4.3.
If we impose the condition of continuity on $T:X\to Y$ we can waive the completeness of $(Y,d_{Y})$.
@@
$J_{M}(t;u,v)=\max\{d_{X}(u,v),\,t\,d_{X}(u,v)\}=\begin{cases}d_{X}(u,v),&0<t\leq 1,\\[4.0pt] t\,d_{X}(u,v),&t\geq 1.\end{cases}$
@@
$d_{X}(u,v)\leq M_{\theta,q}\,p_{\theta,q}(u,v)\qquad\text{and}\qquad p_{\theta,q}(u,v)\leq d_{X}(u,v).$
@@
$\frac{1}{M_{\theta,q}}\,d_{X}(u,v)\;\leq\;p_{\theta,q}(u,v)\;\leq\;d_{X}(u,v).$ (23)
@@
$\frac{1}{M_{\theta,q}}\,d_{X}(u,v)\leq P_{\theta,q}(u,v)\leq d_{X}(u,v),\qquad\forall u,v\in X.$
@@
$\frac{1}{M_{\theta,q}}\,d_{X}(x,y)\leq\delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ inX.$
Therefore $\delta_{\theta,q}$ is equivalent to $d_{X}$. Since $(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}$ was defined as the relative completion of $(X_{0}\cap X_{1},\delta_{\theta,q})$ in $(X_{0}\cup X_{1},K_{M}(1))$ and here $X_{0}\cap X_{1}=X_{0}\cup X_{1}=X$, it holds that
$(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}=X.$
@@
$\frac{1}{M_{\theta,q}}\,d_{X}(x,y)\leq\Delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ \in X$
@@
then $T_{\theta,q}:=T|_{(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}}:(\overrightarrow{X}_{\theta,q}^{J_{M}})^{\prime}\to F$ is Lipschitzian and compact.
@@
Remark 4.4.
The metric interpolation developed in this work can interact with non-linear structures relevant in analysis and PDEs. A natural example is the space of probabilities $\mathcal{P}_{2}(\Omega)$ equipped with the Wasserstein distance $W_{2}$, which is a complete and geodesic metric space, but non-linear. In many evolution problems — such as gradient flows associated with energy functionals in the sense of Jordan–Kinderlehrer–Otto — regularity on different metric scales plays a crucial role. The simultaneous existence of two metric controls (one more “fine”, the other more “coarse”) suggests the possibility of applying metric interpolation constructions to obtain intermediate estimates of stability or temporal continuity.
Although such applications exceed the scope of this article, the abstract structure developed here provides a framework that can be used to investigate gradient flows, metric compactifications, and regularization properties in Wasserstein-type spaces.

### v2

# A new interpolation method for metric spaces based on bi-infinite sequences: The $R$-Method

@@ Abstract
We introduce a new interpolation method for metric spaces, termed the $R$-method, based on bi-infinite linking sequences. Although the construction is inspired by the classical metric functional $J_{M}$, the resulting interpolated space is generated by a distinct object that behaves as a multiscale energy functional. This functional measures the minimal discrete action required to connect two points through $\mathbb{Z}$-indexed sequences, leading to a new intrinsic metric on $X_{0}\cap X_{1}$.
The associated interpolated space is obtained as the relative completion of this metric inside $X_{0}\cup X_{1}$ and is genuinely different from those produced by the $J_{M}$- and $K_{M}$-methods. A fundamental structural property of the $R$-method is that the resulting space embeds continuously into the corresponding $K_{M}$-interpolated space, situating the construction naturally within the existing theory of metric interpolation.
When the method is restricted to a normed setting, the $R$-method induces a genuine interpolation functor. In this framework, it preserves the Lipschitz property of operators with closed graphs, even in the absence of linearity, thereby extending the classical scope of interpolation theory, which is traditionally confined to linear continuous operators. As a consequence, standard compactness properties are also preserved under mild assumptions.
The $R$-method thus provides a new interpolation framework whose foundations rely exclusively on intrinsic metric properties and the summability of discrete orbits, bridging metric interpolation, nonlinear analysis, and classical interpolation theory.
Keywords:
Metric Interpolation. Lipschitz Operators. Nonlinear Analysis. J-method. K-method. Functional Analysis. Interpolation Theory.
Classification codes:
46B70, 54E35, 47H09.
@@ Introduction
The classical theory of interpolation, initiated in the 1960s through the foundational work of Lions and Peetre [10] and later systematized in texts such as [3, 4], is intrinsically rooted in linear and normed structures. Interpolation functors were designed to construct intermediate Banach spaces between two given endpoints, providing powerful tools for the analysis of linear operators, partial differential equations, harmonic analysis, and geometric functional analysis. For decades, the theory evolved almost exclusively within linear frameworks, with its principal constructions relying on convexity, homogeneity, and duality.
Only recently has the need for interpolation tools beyond the linear setting become apparent. Several modern areas, including optimal transport, metric geometry, gradient flows in Wasserstein spaces, and nonlinear analysis on general metric measure spaces, are naturally formulated in purely metric terms and lack vector space structure. Nevertheless, many of these contexts exhibit intermediate regularity phenomena that are conceptually analogous to interpolation. This has motivated the search for interpolation frameworks that operate directly at the level of metric spaces, without recourse to linear or normed structures. Despite progress in nonlinear functional analysis and metric geometry, a systematic and flexible interpolation theory for arbitrary metric spaces has remained largely undeveloped.
In [13], we initiated such a program by introducing two metric-space analogues of Peetre’s classical $K$- and $J$-methods, denoted respectively by the $K_{M}$- and $J_{M}$-methods. These constructions reinterpret the classical ideas entirely in metric terms. Given a compatible pair of metric spaces $(X_{0},X_{1})_{X}$, the $K_{M}$-method is based on admissible finite linking sequences in $X_{0}\cup X_{1}$, penalizing each step through a functional of the form
$K_{M}(t;x,y)=\inf\sum h_{t}(x_{k},x_{k+1}),$
which mirrors the behavior of the classical $K$-functional. In contrast, the $J_{M}$-method acts on the intersection $X_{0}\cap X_{1}$ and is governed by the metric
$J_{M}(t;x,y)=\max\{d_{0}(x,y),\,t\,d_{1}(x,y)\},$
the natural metric counterpart of Peetre’s $J$-functional.
Both methods generate interpolated metric spaces via relative completion: one first endows $X_{0}\cap X_{1}$ with a suitable multiscale metric functional and then completes it inside the ambient space $(X_{0}\cup X_{1},K_{M}(1))$. These constructions preserve Lipschitz continuity of operators, exhibit functorial behavior analogous to that of classical interpolation methods, and apply naturally in purely metric settings, including Wasserstein spaces [1, 15, 16]. Their introduction represents one of the first systematic attempts to extend interpolation theory beyond the linear world.
The present article introduces a third interpolation method, which is structurally distinct from both the $K_{M}$- and $J_{M}$-methods. Although the construction still involves the classical functional $J_{M}$, the underlying mechanism is fundamentally different. Rather than relying on finite admissible chains or single-step comparisons, the new method is built upon bi-infinite linking sequences
$(S_{k})_{k\in\mathbb{Z}}\subset X_{0}\cap X_{1},$
which simultaneously approximate a point $x$ as $k\to+\infty$ and a point $y$ as $k\to$-$\infty$. Along such sequences, we evaluate a multiscale functional of the form
$\Gamma_{\theta,q}\bigl(J_{M}(2^{k};S_{k},S_{k+1})\bigr),$
measuring the minimal discrete action required to connect $x$ and $y$ across all dyadic scales. Taking the infimum over all admissible approximating sequences yields a pre-metric $p_{\theta,q}$; symmetrization and path-length completion then produce a genuine metric $\delta_{\theta,q}$ on $X_{0}\cap X_{1}$.
The resulting interpolated space is defined as the relative completion of $(X_{0}\cap X_{1},\delta_{\theta,q})$ inside $(X_{0}\cup X_{1},K_{M}(1))$. This construction yields a new metric interpolation space, denoted by $R_{\theta,q}^{\overrightarrow{X}}$, whose geometry differs essentially from that of the previously studied $K_{M}$- and $J_{M}$-interpolated spaces. The use of bi-infinite approximating sequences introduces a variational and multiscale structure, capturing information distributed across all dyadic scales. In this respect, the method is closer in spirit to energetic approaches common in modern metric analysis, while remaining firmly rooted in interpolation theory.
From a broader perspective, the development of interpolation theory for metric spaces remains in its early stages, and the availability of multiple mechanisms is essential. Different analytic and geometric problems may require different ways of measuring intermediate regularity or oscillation. The $R$-method enriches the existing toolkit by providing a sequence-based, variational, and genuinely multiscale interpolation mechanism, which may be advantageous in nonlinear settings where convergence of sequences plays a central role, such as gradient flows, geometric evolution problems, and approximation schemes in metric geometry.
The purpose of this article is to present the full construction of the $R$-method, establish its fundamental metric properties, and show that it is compatible with the interpolation of Lipschitz and compact operators. While the method is intrinsically metric, we also show that, when restricted to normed spaces, it induces a genuine interpolation functor, extending the classical scope of interpolation beyond linear operators. In doing so, the article lays the foundations for a new and distinct interpolation mechanism within the emerging theory of metric-space interpolation.
@@ 1.1 Relative Completion of Metric Spaces
We define the relative completion of $(A,d_{A})$ in $(B,d_{B})$, denoted by $A_{B}$, as the set of all the $x\in B$ such that there exists a Cauchy sequence $(x_{n})$ in $(A,d_{A})$ satisfying $d_{B}(x_{n},x)\to 0$.
@@ 1.2 Pairs of Compatible Metric Spaces
Let $(X_{0},X_{1})_{X}$ and $t>0$. For $(x,y)\in X_{0}^{2}\cup X_{1}^{2}$, define the functional hop, denoted by $h_{t}$ which will be given by
@@ 1.3 The $K_{M}$ Functional and its properties
$K_{M}(a;x,y)\leq\max\biggr\{1,\frac{a}{b}\biggr\}\cdot K_{M}(b;x,y)$ (3)
@@ 3 The $J_{M}^{\prime}$-Interpolated Metric Space
Now, using Lemma 5, with $a=1$ and $b=2^{j}$ in each term $K_{M}(1;S_{j}^{x,y},S_{j+1}^{x,y})$, we have
@@
Definition 3.4 (The $R$-interpolated metric space).
The $R$-interpolated metric space is denoted by $R_{\theta,q}^{\overrightarrow{X}}$ and is the relative completion of the metric space
@@
Remark 3.
An important fact is that the interpolated spaces above were defined as the completion of the intersection $X_{0}\cap X_{1}$ with a specific metric in the metric space $(X_{0}\cup X_{1},K_{M}(1))$. This ensures that the elements of the interpolated spaces are contained in $X_{0}\cup X_{1}$. However, note that such metrics on the intersection also satisfy the conditions for completion in $(X,d_{X})$ which can be substantially “larger” than $X_{0}\cup X_{1}$, providing many more new elements to the completions. That is why the name “relative” was chosen for this type of completion. We do not study here the advantages of completing the spaces beyond $X_{0}\cup X_{1}$.
4 Properties of the $R$-Interpolated Metric Space
@@
The inclusions $X_{0}\cap X_{1}\subseteq R_{\theta,q}^{\overrightarrow{X}}\subseteq X_{0}\cup X_{1}$ are valid and furthermore
@@
    $K_{M}(1;x,y)\leq M_{\theta,q}\cdot\Delta_{\theta,q}(x,y),\quad x,y\in R_{\theta,q}^{\overrightarrow{X}}$.
@@
Lemma 9 (Reindexing and interpolation factor).
@@
Then, $T_{\theta,q}:=T|_{R_{\theta,q}^{\overrightarrow{X}}}:R_{\theta,q}^{\overrightarrow{X}}\to R_{\theta,q}^{\overrightarrow{Y}}$ is Lipschitzian with Lipschitz constant not exceeding $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}>0$.
@@
where in the last inequality we used the previous Lemma 9. From this inequality, we have that
@@
$\Delta_{\theta,q;Y}(T(x),T(y))\leq 2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in R_{\theta,q}^{\overrightarrow{X}}$ (19)
@@
$\Delta_{\theta,q;Y}(T(x),T(y))\leq\max\{\omega_{0},\omega_{1}\}\cdot\Delta_{\theta,q;X}(x,y),\quad x,y\in R_{\theta,q}^{\overrightarrow{X}}.$ (20)
@@
Remark 6.
In Theorem 4.2 if we impose the condition of continuity on $T:X\to Y$ we can waive the completeness of $(Y,d_{Y})$.
@@
$J_{M}(t;u,v)=\max\{d_{X}(u,v),\,t\cdot d_{X}(u,v)\}=\begin{cases}d_{X}(u,v),&0<t\leq 1,\\ t\cdot d_{X}(u,v),&t\geq 1.\end{cases}$
@@
$d_{X}(u,v)\leq M_{\theta,q}\cdot p_{\theta,q}(u,v)\qquad\text{and}\qquad p_{\theta,q}(u,v)\leq d_{X}(u,v).$
@@
$\frac{1}{M_{\theta,q}}\cdot d_{X}(u,v)\;\leq\;p_{\theta,q}(u,v)\;\leq\;d_{X}(u,v).$ (23)
@@
$\frac{1}{M_{\theta,q}}\cdot d_{X}(u,v)\leq P_{\theta,q}(u,v)\leq d_{X}(u,v),\qquad\forall u,v\in X.$
@@
$\frac{1}{M_{\theta,q}}\cdot d_{X}(x,y)\leq\delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ inX.$
Therefore $\delta_{\theta,q}$ is equivalent to $d_{X}$. Since $R_{\theta,q}^{\overrightarrow{X}}$ was defined as the relative completion of $(X_{0}\cap X_{1},\delta_{\theta,q})$ in $(X_{0}\cup X_{1},K_{M}(1))$ and here $X_{0}\cap X_{1}=X_{0}\cup X_{1}=X$, it holds that
$R_{\theta,q}^{\overrightarrow{X}}=X.$
@@
$\frac{1}{M_{\theta,q}}\cdot d_{X}(x,y)\leq\Delta_{\theta,q}(x,y)\leq d_{X}(x,y),\qquad\forall x,y\ \in X$
@@
then $T_{\theta,q}:=T|_{R_{\theta,q}^{\overrightarrow{X}}}:R_{\theta,q}^{\overrightarrow{X}}\to F$ is Lipschitzian and compact.
@@
Corollary 3.
Let $(X_{0},X_{1})_{X}$ be a pair of compatible metric spaces, $\theta\in(0,1)$, $q\in[1,\infty)$ and $T:X\to X$ a closed operator that satisfies
  • (a)
    $T_{0}:X_{0}\rightarrow X_{0}$ is a lipschitzian operator with constant $\omega_{0}$;
  • (b)
    $T_{1}:X_{1}\rightarrow X_{1}$ is a lipschitzian operator with constant $\omega_{1}$;
  • (c)
    $X_{1}$ is closed in $(X_{0}\cup X_{1},K_{M}(1))$ and $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}<1$.
Then $T_{1}$ has a fixed point.
Proof.
Due to Theorem 4.2, the operator $T$ restricted to the interpolated space $R_{\theta,q}^{\overrightarrow{X}}$ is lipschitzian with constant not exceeding $2^{\theta}\cdot\omega_{0}^{1-\theta}\cdot\omega_{1}^{\theta}$. Thus, $T_{\theta,q}$ is a strong contraction. Since $(X,d_{X})$ is complete, it follows that the relative completion is also complete. Thus, by Banach’s Fixed-Point Theorem, the operator $T_{\theta,q}$ has a unique fixed point $w$ in $R_{\theta,q}^{\overrightarrow{X}}$. By the density of $X_{0}\cap X_{1}$ in the interpolated space and the continuous inclusion of the interpolated space in the metric space $(X_{0}\cup X_{1},K_{M}(1))$, there exists a sequence $(w_{n})$ of points in $X_{0}\cap X_{1}$ such that $K_{M}(1;w_{n},w)\to 0$. Since $X_{1}$ is closed in $(X_{0}\cup X_{1},K_{M}(1))$, it follows that $w\in X_{1}$. Thus, due to the closedness of $T$ in all $X$, it follows that $T_{1}(w)=w$.
5 The $R$-method in Banach Spaces
Let $(X_{0},\|\cdot\|_{0})$ and $(X_{1},\|\cdot\|_{1})$ be Banach spaces continuously embedded into a common Hausdorff topological vector space $Z$ over a field $\mathbb{F}$ ($\mathbb{R}$ or $\mathbb{C}$), with $X_{0}\cap X_{1}\neq\emptyset$. Let us consider the classic $K$-functional due to Lions-Peetre[see [10]] given by
$K(t;x):=\inf_{x_{0}+x_{1}=x}(\|x_{0}\|_{0}+t\cdot\|x_{1}\|_{1}),\quad t>0,\quad x\in X_{0}+X_{1}.$ (24)
It is known that for all $t>0$ the functional $K(t;\cdot)$ is a norm on $X=X_{0}+X_{1}$ and that $(X_{0}+X_{1},K(t;\cdot))$ is a Banach space.
Note that $K(1;x)\leq\|x\|_{0}$ for all $x\in X_{0}$ and $K(1;y)\leq\|y\|_{1}$ for all $y\in X_{1}$. It shows that $(X_{0},X_{1})_{X}$ is a compatible pair of metric spaces. In this environment, we have
Proposition 5.1.
$R_{\theta,q}^{\overrightarrow{X}}$ is a Banach space.
Proof.
Note that $0\in X_{0}\cap X_{1}\subseteq R_{\theta,q}^{\overrightarrow{X}}$. Now, we define $\|x\|_{\theta,q}:=\Delta_{\theta,q}(x,0)$ for all $x\in R_{\theta,q}^{\overrightarrow{X}}$. The set $X_{0}\cap X_{1}$ is endowed with a linear structure. By density of $X_{0}\cap X_{1}$ we can define a linear structure in $R_{\theta,q}^{\overrightarrow{X}}$ by putting for all $x,y\in R_{\theta,q}^{\overrightarrow{X}}$ and $\lambda\in\mathbb{F}$
$x+\lambda\cdot y:=\lim_{n\to\infty}x_{n}+\lambda\cdot y_{n},$ (25)
where $\|x_{n}-x\|_{\theta,q}\to 0$ and $\|y_{n}-y\|_{\theta,q}\to 0$ and $(x_{n})$ and $(y_{n})$ are sequences of points in $X_{0}\cap X_{1}$ given by the density. By the equality 25, it follows that $\|\cdot\|_{\theta,q}$ is a norm on $R_{\theta,q}^{\overrightarrow{X}}$ because $\delta_{\theta,q}=\Delta_{\theta,q}$ in $X_{0}\cap X_{1}$. Furthermore, since $(X_{0},\|\cdot\|_{0})$ and $(X_{1},\|\cdot\|_{1})$ are Banach spaces, it is also known that $X=X_{0}+X_{1}$ is a Banach space with the norm $K(1;\cdot)$. Thus, $R_{\theta,q}^{\overrightarrow{X}}$ is a complete metric space, because it is the relative completion of $(X_{0}\cap X_{1},\|\cdot\|_{\theta,q})$ in $(X_{0}+X_{1},K(1;\cdot))$ and therefore a Banach Space.
Remark 7.
This feature highlights a genuine advantage of the $R$-method. Classical interpolation functors on normed spaces are formulated within a linear framework and act exclusively on linear continuous operators. By contrast, the $R$-method is intrinsically metric and does not rely on linear structure. When restricted to normed spaces, it still induces an interpolation functor and preserves the Lipschitz property of operators with closed graphs, allowing for the interpolation of nonlinear mappings.
In particular, the method enlarges the classical category of admissible operators while retaining the functorial behavior expected in interpolation theory, thereby providing a natural extension of classical interpolation beyond the linear setting.